What is the sum of the geometric sequence 1, 4, 16, ... if there are 8 terms?
Answer by stanbon(75887)

What is the sum of the geometric sequence 1, 4, 16, ... if there are 8 terms?
--------
n = 8 ; a = 1 ; r = 4/1 = 4
--------
S(n) = a[(r^n-1]/[r-1]
S(8) = 1[255/3] = 85
=========================
Cheers,
Stan H.
=================